Calculate the angle between r and positive z axis. Find the angle between a and b. Where a=5i+4j-6k,b=-2i+2j+4k,r=a+b+c. 

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
vector c is not given hence I am only able to soleve nagle between a and b only.
 
- 10 + 8 – 24 = sqrt(77) sqrt (24) cos theta
 
hence cos theta = – 26/ sqrt(77) sqrt(24)
Vikas TU
4 years ago
Dear student 
Please provide the vector C 
Because r = a+b+c 
If we do not know C how can we calculate C , 
Later on the r can not be calculted . 
So, provide the whole question .